OpenStudy (anonymous):

What is the ratio 16 to 40 written as a fraction in lowest terms? A. 1/4 B. 8/20 Ratio, Proportion, and Percent C. 2/5 Ratio, Proportion, and Percent D. 3/8 Ratio, Proportion, and Percent

OpenStudy (welshfella):

you need to divide top and bottom of the fraction by the GCF (reates common factor) of 16 and 40. 16 = 2*2*2*2 40= 2*2*2*5

OpenStudy (welshfella):

pick out the factors that are in both lists

OpenStudy (welshfella):

* greatest common factor